OpenStudy (anonymous):

Will give a medal! A line has a slope of -3 and passes through the point (5, -1). What is the equation of the line? My guess is -3x + 2 = -2

OpenStudy (cooltowl):

yes I believe the equation you use is y=mx+b

OpenStudy (onepieceftw):

You use point-slope form y-y_1 = m(x-x_1)

OpenStudy (anonymous):

ok so it is something y - x = 14

OpenStudy (onepieceftw):

Nope. y-1=-3(x-5) is it in point-slope form Distribute what's in the parentheses to get y-1=-3+15 add 1 to both sides y=-3x+16